OpenStudy (anonymous):

Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5). -6 6 8

OpenStudy (anonymous):

Substitute and calculate. It doesn't get much more straightforward.

OpenStudy (anonymous):

so start with the -2 and plug it in for X and then do it again with 5?

hero (hero):

Hint: f(-2) = 3(-2) + 1 g(5) = -(5) + 6

OpenStudy (texaschic101):

you figure it out yet ?

OpenStudy (anonymous):

i think so

OpenStudy (texaschic101):

what did you get ?

OpenStudy (anonymous):

6

OpenStudy (texaschic101):

f(x) = 3x - 1 f(-2) = 3(-2) - 1 f(-2) = -6 - 1 f(-2) = -7 g(x) = -x + 6 g(5) = -5 + 6 g(5) = 1 f(x) + g(x) = -7 + 1 = -6

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

yes thank you

OpenStudy (texaschic101):

no problem :) You got to be careful with the signs

OpenStudy (anonymous):

kinda sucks but yeah haha time for the test _-_

OpenStudy (texaschic101):

good luck :)

OpenStudy (anonymous):

thanks